OpenStudy (anonymous):

Find the value of cos^2 45- Sin^2 15 Ans is sqrt(3)/4

OpenStudy (anonymous):

\[\cos ^2 45- \sin^2 15= ? \]

OpenStudy (anonymous):

cos^a -sin^b= cos(a+b)*cos(a-b)

OpenStudy (anonymous):

use the relation cos^2a -sin^2b= cos(a+b)*cos(a-b)

OpenStudy (anonymous):

cos^2 45- Sin^2 15 =cos(45+15)*cos(45-15)=cos60 *cos30=1/2 *sqrt(3) /2= sqrt(3) /4

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

welcome